# Corresponding author(s)

### v1

@@ 4.2 Main Results
Figure 5: Comparative results for different difficulty levels on MATH500
@@
4.3 Case Study
Figure 6 and 7 demonstrate a comparison of results for two simple problems from DeepSeek-R1-Distill-Qwen-32B. It can be observed that that the original outputs include extensive unnecessary and redundant thinking processes, while the outputs after applying DAST are more concise and focused. Figure 8 demonstrates that the DAST method maintains its reasoning capability in addressing complex problems.
Figure 8: A Level 5 case from MATH500 answered by DeepSeek-R1-Distill-Qwen-32B with DAST.
5 Conclusion
This work addresses the critical efficiency-performance dilemma in slow-thinking models through difficulty-aware reasoning adaptation. By establishing correlation between problem complexity and optimal solution length, the proposed DAST framework enables dynamic resource allocation for reasoning. Experimental validations across MATH500 and AIME2024 confirm the effectiveness of our method.

### v2

@@ 4.2.1 Overall Results
Figure 5: Comparative results for different difficulty levels on MATH-500
4.2.2 Fine-grained Analysis
We also compared CR of DAST (DS-32B) on MATH-500 according to difficulty level. As shown in Figures 5, DAST achieved the best Level 5 accuracy with significant margin against other methods, which demonstrates that it maintains its reasoning ability under complex problems. Although SimPOShortest shows the most significant reduction in response length, its reasoning capability notably declines when addressing complex problems.
Model L1 L2 L3 L4 L5
SimPOShortest 72.0% 64.1% 67.6% 72.1% 72.5%
DAST 58.5% 47.7% 51.9% 46.7% 40.8%
Table 2: Comparison of compression ratio between SimPOShortest and DAST across different levels in MATH-500 with DS-32B.
Statistical analysis of compression ration at different levels on MATH-500 in Table 2 reveals that SimPOShortest exhibits limited differentiation in CR across different difficulty levels. In contrast, the DAST method shows discernible adaptive capabilities, achieving approximately 58.5% length reduction at Level 1 compared to the original model, while this reduction decreases to approximately 40.8% at the most challenging Level 5. This progressive performance variation validating its difficulty-adaptive nature.
4.2.3 Ablation Study
To reveal the individual effects of different components of our method, we tested different variants of DAST on MATH-500 with DS-7B by removing DCP or DICP. The ablation results are shown in Table 3. We see that the DCP and DICP components exhibit specialization patterns analogous to domain-specific experts. When removing DCP while retaining DICP (w/o DCP), the framework incentivizes models to fully utilize the token budget, resulting in an accuracy improvement(+1.4% versus DS-7B). However, this comes at the cost of overly redundant answer length (+17.8% versus DS-7B). Conversely, eliminating DICP while preserving DCP (w/o DCIP) drives the model to strictly adhere to budget constraints through aggressive compression, achieving optimal compression ratio (59.8%) but significantly impairing problem-solving capability (-3.2% accuracy). The optimal performance is achieved when DCP and DICP are combined, indicating that DCP and DICP are complementary to each other. We further explored integrating CICP into DAST’s training set (incorporating CICPs with the largest reward score discrepancies per question). However, as evidenced in Table 3 (bottom row), this integration yielded no significant performance gains. We will investigate the ineffectiveness of CICP in the future work.
Model ACC LEN C-LEN CR
DS-7B 93.2 4039.13 3506.64 -
DAST 93.6 3309.16 2708.63 18.0%
w/o DCP 94.6 4759.07 3996.78 -17.8%
w/o DICP 90.0 1624.60 1299.50 59.8%
+ CICP 93.2 3295.96 2573.00 18.3%
Table 3: Ablation Results on MATH-500 with DS-7B.
4.2.4 Impact of Truncation Threshold
To investigate the impact of the truncation threshold $\delta$, we conducted grid search validation on 100 randomly selected samples from MATH-500. As shown in Figure 6, the DS-32B model achieves peak ACC with $\delta$ = 0.15, accompanied by a CR of 47% in generated token length. This empirical evidence guided our final selection of $\delta$ = 0.15 for DS-32B to optimize the Accuracy. For the DS-7B variant, the same hyperparameter search on the same validation set identified 0.18 as the optimal $\delta$.
It is noteworthy that when $\delta$ = 0, the model’s CR becomes extremely low (even negative), primarily because the training data contains DICPs with low discriminability and excessive length, which causes reward hacking and prevents SimPO from capturing the correct direction for length optimization.
Figure 6: The impact of truncation threshold $\delta$.
5 Conclusion
This work addresses the critical efficiency-performance dilemma in slow thinking models through difficulty-aware reasoning adaptation. By establishing correlation between problem complexity and optimal solution length, the proposed DAST framework enables dynamic resource allocation for reasoning. Experimental validations across representative benchmarks confirm the effectiveness of our method.
Limitations
While our introduced method achieves a remarkable trade-off between reasoning accuracy and response compression rate, following limitations warrant discussion:
Domain-Specific Evaluation Scope
Our current benchmarking focuses exclusively on STEM disciplines (e.g., mathematics, physics, chemistry), leaving code generation and general domain tasks unexplored. We plan to extend the evaluation benchmarks in the future.
Threshold Sensitivity
Our method is sensitive to the truncation threshold. Therefore, it requires some additional cost to carefully adjust the threshold.
Off-Policy Learning Constraints
The proposed DAST framework, though computationally efficient through preconstructed training data, may inherently limit performance potential compared to online reinforcement learning approaches. We plan to explore on-policy reinforcement learning variants using our designed reward function for further improvement.
Appendix A Implementation Details
Model Name Value
DS-7B training samples 10295
learning rate 5e-6
DCP % 91.26%
DCIP % 8.74%
epoch 1
$L_{max}$ 4096
$\beta$ 200
$\gamma$ 1
DS-32B training samples 9813
learning rate 5e-6
DCP % 87.73%
DCIP % 12.27%
epoch 1
$L_{max}$ 4096
$\beta$ 200
$\gamma$ 1
Table 4: Training configuration of DAST.
Appendix B Prompt Templates of CCoT and CoD
Figure 7: The prompt we used to implement CCoT method.
Figure 8: The prompt we used to implement CoD method.
Appendix C Case Study
Figure 9 demonstrate a comparison of results for a simple problem from DeepSeek-R1-Distill-Qwen-32B. It can be observed that that the original outputs include extensive unnecessary and redundant thinking processes, while the outputs after applying DAST are more concise and focused.